6) A 45 ml 0.09 M KOH solution is titrated with 0.11 M HCI: a) How many mL of HCI solution are required to reach the equivalence point? b) What is the pH at the midpoint of the titration? c) What is the pH at the equivalence point?
Answer #1

a)

At equivalence point, number of moles of HCl will be equal to number of moles of KOH

HCl + KOH ----------- KCl + H2O

volume of KOH (in mL) * molarity of KOH = volume of HCl (in mL) * molarity of HCl

45 * 0.09 = V * 0.11

V = 36.82 mL

b) Half mid point of titration

Volume of HCl used = 36.82/2 = 18.41 mL

Total volume = 45 + 18.41 = 63.41 mL

Number of moles of KOH left = 1/2 * 45/1000 * 0.09 = 0.002025 moles

Molarity of KOH = 0.002025/63.41 * 1000 = 0.0319 M

pOH = -log[OH-] = -log[0.0319] = 1.49

pH = 14 - pOH = 14 - 1.49 = 12.51

c) pH at equivalence point will be equal to 7 since the base will be completely neutralized by the acid